# Classification of nilpotent almost abelian Lie groups admitting left-invariant complex or symplectic structures

### v1

@@ 2. ${\rm(ii)}$
Let $\mathfrak{g}={\mathbb{R}}\oplus\mathfrak{h}_{5}$, where $\mathfrak{h}_{5}$ is the $5$-dimensional Heisenberg Lie algebra. Then $\mathfrak{g}$ admits a complex structure but no symplectic structure (see [38] and [25], respectively). This Lie algebra was denoted by $(0,0,0,0,0,12+34)$ in [38].
5. Applications
5.1. Low dimensions
In this section, we apply Theorems 3.10 and 4.10 to study which nilpotent almost abelian Lie algebras of low dimensions admit complex or symplectic structures. We focus our attention on dimensions $8$ and $10$ since the existence of complex and symplectic structures on $6$-dimensional nilpotent Lie algebras has been already studied in [38] and [25], respectively.
In the following two corollaries, we study the $8$-dimensional case.
Corollary 5.1.
Let $\mathfrak{g}_{A}$ be a nilpotent almost abelian Lie algebra of dimension $8$. Then, $\mathfrak{g}_{A}$ admits a complex structure if and only if $A$ is conjugate to one of the following matrices:
$\mathcal{J}_{4}\oplus\mathcal{J}_{3},\;\ \mathcal{J}_{3}^{\oplus 2}\oplus 0_{1},\;\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}\oplus 0_{2},\;\ \mathcal{J}_{2}^{\oplus 3}\oplus 0_{1},\;\ \mathcal{J}_{2}^{\oplus 2}\oplus 0_{3},\;\ \mathcal{J}_{2}\oplus 0_{5}.$
Proof.
Since $\mathfrak{g}_{A}$ is a nilpotent almost abelian Lie algebra, $A$ is a nilpotent $7\times 7$ matrix. Therefore, by the Jordan normal form, $A$ will be conjugated to one of the following matrices (according to its step of nilpotency):
(12) $\begin{array}[]{cccccc}\mathcal{J}_{7};&\mathcal{J}_{6}\oplus 0_{1};&\begin{cases}\mathcal{J}_{5}\oplus\mathcal{J}_{2}\\ \mathcal{J}_{5}\oplus 0_{2}\end{cases}\!\!\!\!\!\!;&\begin{cases}\mathcal{J}_{4}\oplus\mathcal{J}_{3}\\ \mathcal{J}_{4}\oplus\mathcal{J}_{2}\oplus 0_{1}\\ \mathcal{J}_{4}\oplus 0_{3}\end{cases}\!\!\!\!\!\!;&\begin{cases}\mathcal{J}_{3}^{\oplus 2}\oplus 0_{1}\\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}^{\oplus 2}\\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}\oplus 0_{2}\\ \mathcal{J}_{3}\oplus 0_{4}\end{cases}\!\!\!\!\!\!;&\begin{cases}\mathcal{J}_{2}^{\oplus 3}\oplus 0_{1}\\ \mathcal{J}_{2}^{\oplus 2}\oplus 0_{3}\\ \mathcal{J}_{2}\oplus 0_{5}\end{cases}\!\!\!\!\!\!.\end{array}$
It is easy to check that the matrices in (12) satisfying (i), (ii) or (iii) in Theorem 3.10 are
$\mathcal{J}_{4}\oplus\mathcal{J}_{3},\;\ \mathcal{J}_{3}^{\oplus 2}\oplus 0_{1},\;\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}\oplus 0_{2},\;\ \mathcal{J}_{2}^{\oplus 3}\oplus 0_{1},\;\ \mathcal{J}_{2}^{\oplus 2}\oplus 0_{3},\;\ \mathcal{J}_{2}\oplus 0_{5}.\qed$
In contrast with the complex case, it follows that any $8$-dimensional nilpotent almost abelian Lie algebra admits a symplectic structure, as the next result shows.
Corollary 5.2.
Let $\mathfrak{g}_{A}$ be a nilpotent almost abelian Lie algebra of dimension $8$. Then, $\mathfrak{g}_{A}$ admits a symplectic structure.
Proof.
It is easy to check that any matrix in (12) satisfies (i) or (ii) in Theorem 4.10. ∎
Recall that we obtained in Example 4.7 a $10$-dimensional $5$-step nilpotent Lie algebra admitting no symplectic structure. According to Theorem 4.6 and Corollary 5.2 this is the lowest degree of nilpotency and the lowest dimension where this occurs.
Now, we study the $10$ dimensional case.
Corollary 5.3.
Let $\mathfrak{g}_{A}$ be a nilpotent almost abelian Lie algebra of dimension $10$. Then, $\mathfrak{g}_{A}$ admits a complex structure if and only if $A$ is conjugate to one of the following matrices:
$\mathcal{J}_{5}\oplus\mathcal{J}_{4},\;\ \mathcal{J}_{4}^{\oplus 2}\oplus 0_{1},$ $\;\ \mathcal{J}_{4}\oplus\mathcal{J}_{3}\oplus 0_{2},\;\ \mathcal{J}_{3}^{\oplus 2}\oplus\mathcal{J}_{2}\oplus 0_{1},\;\ \mathcal{J}_{3}^{\oplus 2}\oplus 0_{3},\;\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}^{\oplus 3},$
$\mathcal{J}_{3}\oplus\mathcal{J}_{2}\oplus 0_{4},$ $\;\ \mathcal{J}_{2}^{\oplus 4}\oplus 0_{1},\;\ \mathcal{J}_{2}^{\oplus 3}\oplus 0_{3},\;\ \mathcal{J}_{2}^{\oplus 2}\oplus 0_{5},\;\ \mathcal{J}_{2}\oplus 0_{7}.$
Proof.
Since $\mathfrak{g}_{A}$ is a nilpotent almost abelian Lie algebra, $A$ is a nilpotent $9\times 9$ matrix. Therefore, by the Jordan normal form, $A$ will be conjugated to one of the following matrices (according to its step of nilpotency):
(13) $\begin{array}[]{llll}\mathcal{J}_{9};&\mathcal{J}_{8}\oplus 0_{1};&\begin{cases}\mathcal{J}_{7}\oplus\mathcal{J}_{2}\\ \mathcal{J}_{7}\oplus 0_{2}\end{cases}\!\!\!\!\!;&\begin{cases}\mathcal{J}_{6}\oplus\mathcal{J}_{3}\\ \mathcal{J}_{6}\oplus\mathcal{J}_{2}\oplus 0_{1}\\ \mathcal{J}_{6}\oplus 0_{3}\end{cases}\!\!\!\!\!\!;\\ \begin{cases}\mathcal{J}_{5}\oplus\mathcal{J}_{4}\\ \mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}\\ \mathcal{J}_{5}\oplus\mathcal{J}_{2}^{\oplus 2}\\ \mathcal{J}_{5}\oplus\mathcal{J}_{2}\oplus 0_{2}\\ \mathcal{J}_{5}\oplus 0_{4}\end{cases}\!\!\!\!\!\!;&\begin{cases}\mathcal{J}_{4}^{\oplus 2}\oplus 0_{1}\\ \mathcal{J}_{4}\oplus\mathcal{J}_{3}\oplus\mathcal{J}_{2}\\ \mathcal{J}_{4}\oplus\mathcal{J}_{3}\oplus 0_{2}\\ \mathcal{J}_{4}\oplus\mathcal{J}_{2}^{\oplus 2}\oplus 0_{1}\\ \mathcal{J}_{4}\oplus\mathcal{J}_{2}\oplus 0_{3}\\ \mathcal{J}_{4}\oplus 0_{5}\end{cases}\!\!\!\!\!\!;&\begin{cases}\mathcal{J}_{3}^{\oplus 3}\\ \mathcal{J}_{3}^{\oplus 2}\oplus\mathcal{J}_{2}\oplus 0_{1}\\ \mathcal{J}_{3}^{\oplus 2}\oplus 0_{3}\\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}^{\oplus 3}\\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}^{\oplus 2}\oplus 0_{2}\\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}\oplus 0_{4}\\ \mathcal{J}_{3}\oplus 0_{6}\end{cases}\!\!\!\!\!\!;&\begin{cases}\mathcal{J}_{2}^{\oplus 4}\oplus 0_{1}\\ \mathcal{J}_{2}^{\oplus 3}\oplus 0_{3}\\ \mathcal{J}_{2}^{\oplus 2}\oplus 0_{5}\\ \mathcal{J}_{2}\oplus 0_{7}\end{cases}\!\!\!\!\!\!.\end{array}$
It is easy to check that the above matrices satisfying (i), (ii) or (iii) in Theorem 3.10 are
$\mathcal{J}_{5}\oplus\mathcal{J}_{4},\;\ \mathcal{J}_{4}^{\oplus 2}\oplus 0_{1},$ $\;\ \mathcal{J}_{4}\oplus\mathcal{J}_{3}\oplus 0_{2},\;\ \mathcal{J}_{3}^{\oplus 2}\oplus\mathcal{J}_{2}\oplus 0_{1},\;\ \mathcal{J}_{3}^{\oplus 2}\oplus 0_{3},\;\ \mathcal{J}_{3}\oplus\mathcal{J}_{2}^{\oplus 3},$
$\mathcal{J}_{3}\oplus\mathcal{J}_{2}\oplus 0_{4},$ $\;\ \mathcal{J}_{2}^{\oplus 4}\oplus 0_{1},\;\ \mathcal{J}_{2}^{\oplus 3}\oplus 0_{3},\;\ \mathcal{J}_{2}^{\oplus 2}\oplus 0_{5},\;\ \mathcal{J}_{2}\oplus 0_{7}.\qed$
In Example 4.7 we exhibit the Lie algebra $\mathfrak{g}_{M}$ with $M=\mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}$, which admits no symplectic structure. Analyzing case by case all the square matrices given in (13), it is easy to see that the only one that satisfies neither (i) nor (ii) of Theorem 4.10 is $\mathfrak{g}_{M}$. It follows from Theorem 4.11 that $\mathfrak{g}_{M}$ admits no complex structure.
Corollary 5.4.
Up to isomorphism, the only $10$-dimensional nilpotent almost abelian Lie algebra admitting neither complex nor symplectic structure is $\mathfrak{g}_{M}$ with $M=\mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}$.
We observe that, when $M=\mathcal{J}_{m_{1}}\oplus\mathcal{J}_{m_{2}}\oplus 0_{1}$ with $m_{1},\,m_{2}$ odd, $m_{1}>m_{2}$, then $\mathfrak{g}_{M}$ admits no symplectic structure. Therefore by Theorem 4.11 $\mathfrak{g}_{M}$ admits neither complex nor symplectic structure (see Corollary 5.8 for other families of examples).
5.2. Further consequences of Theorems 3.10 and 4.10
@@ 5.1. Low dimensions
$M=\mathcal{J}_{6}^{\oplus p_{6}}\oplus\mathcal{J}_{5}^{\oplus p_{5}}\oplus\mathcal{J}_{4}^{\oplus p_{4}}\oplus\mathcal{J}_{3}^{\oplus p_{3}}\oplus\mathcal{J}_{2}^{\oplus p_{2}}\oplus 0_{t},\quad p_{6}\geq 1,\;\;t\geq 0,$
with $\sum_{i=2}^{6}ip_{i}+t=2n-1$.
Since $\sum_{i=2}^{6}ip_{i}+t$ is odd , it follows that $5p_{5}+3p_{3}+t$ is odd. If $t$ is even, then $p_{5}+p_{3}$ is odd, so only one of $p_{5}$ and $p_{3}$ is odd, then condition ${\rm(ii)}$ of Theorem 4.10 is satisfied and $\mathfrak{g}_{M}$ admits a symplectic structure. So, by Remark 2.2, $\mathfrak{g}_{A}$ also admits a symplectic structure. If $t$ is odd ($t\geq 1$), then $p_{5}+p_{3}$ is even, so $p_{5}$ and $p_{3}$ are both odd or both even. Suppose $p_{5}$ and $p_{3}$ are both odd ($p_{5}\geq 1$ and $p_{3}\geq 1$), we have
@@
$M=\mathcal{J}_{k}^{\oplus p_{k}}\oplus\cdots\oplus\mathcal{J}_{4}^{\oplus p_{4}}\oplus\mathcal{J}_{3}^{\oplus p_{3}}\oplus\mathcal{J}_{2}^{\oplus p_{2}}\oplus 0_{t},\quad p_{k}\geq 1,\;\;t\geq 0,$
with $\sum_{i=2}^{k}ip_{i}+t=2n-1$.
Since $p_{k}\geq 1$ we have
@@
The upper bound $k+8$ in Corollary 5.6 is optimal, since if we consider even $k$, $k\geq 6$ and $M=\mathcal{J}_{k}\oplus\mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}$, then $\mathfrak{g}_{M}$ is a $k$-step nilpotent almost abelian Lie algebra of dimension $k+10$, but by Theorem 4.10 $\mathfrak{g}_{M}$ does not admit a symplectic structure.
@@
$M=\mathcal{J}_{k}^{\oplus p_{k}}\oplus\cdots\oplus\mathcal{J}_{4}^{\oplus p_{4}}\oplus\mathcal{J}_{3}^{\oplus p_{3}}\oplus\mathcal{J}_{2}^{\oplus p_{2}}\oplus 0_{t},\quad p_{k}\geq 1,\;\;t\geq 0,$
with $\sum_{i=2}^{k}ip_{i}+t=2n-1$.
Since $p_{k}\geq 1$ we have
@@
Then $k(p_{k}-1)+\sum_{i=2}^{k-1}ip_{i}+t$ is even and less than or equal to $2$. Therefore $M=\mathcal{J}_{k}$, $M=\mathcal{J}_{k}\oplus\mathcal{J}_{2}$ or $M=\mathcal{J}_{k}\oplus 0_{2}$, any of the three possibilities admits symplectic structure by Theorem 4.10. Using again Remark 2.2, we obtain that $\mathfrak{g}_{A}$ admits a symplectic structure. ∎
The upper bound $k+3$ in Corollary 5.7 is optimal, since if we consider an odd integer $k$, $k\geq 5$ and $M=\mathcal{J}_{k}\oplus\mathcal{J}_{3}\oplus 0_{1}$, then $\mathfrak{g}_{M}$ is a $k$-step nilpotent almost abelian Lie algebra of dimension $k+5$, but by Theorem 4.10 $\mathfrak{g}_{M}$ does not admit a symplectic structure.
As a final application of Theorems 3.10 and 4.10, we obtain families of nilpotent almost abelian Lie algebras admitting both, complex and symplectic structures, or symplectic but no complex structure, or neither complex nor symplectic structure.

### v2

@@ 2. ${\rm(ii)}$
Let $\mathfrak{g}={\mathbb{R}}\oplus\mathfrak{h}_{5}$, where $\mathfrak{h}_{5}$ is the $5$-dimensional Heisenberg Lie algebra. Then $\mathfrak{g}$ admits a complex structure but no symplectic structure (see [42] and [27], respectively).
5.3. Applications of Theorem 5.10
@@
$M=\mathcal{J}_{6}^{\oplus p_{6}}\oplus\mathcal{J}_{5}^{\oplus p_{5}}\oplus\mathcal{J}_{4}^{\oplus p_{4}}\oplus\mathcal{J}_{3}^{\oplus p_{3}}\oplus\mathcal{J}_{2}^{\oplus p_{2}}\oplus 0_{t},\quad p_{6}\geq 1,\;\;t\geq 0,\quad\sum_{i=2}^{6}ip_{i}+t=2n-1,$
where $p_{i}\geq 0$ for $i=2,\ldots,5$. The meaning of $p_{i}=0$ is that the Jordan block $\mathcal{J}_{i}$ does not appear in $M$. Since $\sum_{i=2}^{6}ip_{i}+t$ is odd, it follows that $5p_{5}+3p_{3}+t$ is odd. If $t$ is even, then $p_{5}+p_{3}$ is odd, so only one of $p_{5}$ and $p_{3}$ is odd, then condition ${\rm(ii)}$ of Theorem 5.10 is satisfied and $\mathfrak{g}_{M}$ admits a symplectic structure. So, by Remark 2.2, $\mathfrak{g}_{A}$ also admits a symplectic structure. If $t$ is odd ($t\geq 1$), then $p_{5}+p_{3}$ is even, so $p_{5}$ and $p_{3}$ are both odd or both even. Suppose $p_{5}$ and $p_{3}$ are both odd ($p_{5}\geq 1$ and $p_{3}\geq 1$), we have
@@
$M=\mathcal{J}_{k}^{\oplus p_{k}}\oplus\cdots\oplus\mathcal{J}_{4}^{\oplus p_{4}}\oplus\mathcal{J}_{3}^{\oplus p_{3}}\oplus\mathcal{J}_{2}^{\oplus p_{2}}\oplus 0_{t},\quad p_{k}\geq 1,\;\;t\geq 0,\quad\sum_{i=2}^{k}ip_{i}+t=2n-1,$
where $p_{i}\geq 0$ for $i=2,\ldots,k-1$. The meaning of $p_{i}=0$ is that the Jordan block $\mathcal{J}_{i}$ does not appear in $M$. Since $p_{k}\geq 1$ we have
@@
The upper bound $k+8$ in Corollary 5.19 is optimal, since if we consider even $k\geq 6$ and $M=\mathcal{J}_{k}\oplus\mathcal{J}_{5}\oplus\mathcal{J}_{3}\oplus 0_{1}$, then $\mathfrak{g}_{M}$ is a $k$-step nilpotent almost abelian Lie algebra of dimension $k+10$, but by Theorem 5.10 $\mathfrak{g}_{M}$ does not admit a symplectic structure.
@@
$M=\mathcal{J}_{k}^{\oplus p_{k}}\oplus\cdots\oplus\mathcal{J}_{4}^{\oplus p_{4}}\oplus\mathcal{J}_{3}^{\oplus p_{3}}\oplus\mathcal{J}_{2}^{\oplus p_{2}}\oplus 0_{t},\quad p_{k}\geq 1,\;\;t\geq 0,\quad\sum_{i=2}^{k}ip_{i}+t=2n-1,$
where $p_{i}\geq 0$ for $i=2,\ldots,k-1$. As before, the meaning of $p_{i}=0$ is that the Jordan block $\mathcal{J}_{i}$ does not appear in $M$. Since $p_{k}\geq 1$ we have
@@
Then $k(p_{k}-1)+\sum_{i=2}^{k-1}ip_{i}+t$ is even and less than or equal to $2$. Therefore, $M=\mathcal{J}_{k}$, $M=\mathcal{J}_{k}\oplus\mathcal{J}_{2}$ or $M=\mathcal{J}_{k}\oplus 0_{2}$, any of the three possibilities admits symplectic structure by Theorem 5.10. Using again Remark 2.2, we obtain that $\mathfrak{g}_{A}$ admits a symplectic structure. ∎
The upper bound $k+3$ in Corollary 5.20 is optimal, since if we consider an odd integer $k\geq 5$ and $M=\mathcal{J}_{k}\oplus\mathcal{J}_{3}\oplus 0_{1}$, then $\mathfrak{g}_{M}$ is a $k$-step nilpotent almost abelian Lie algebra of dimension $k+5$, but by Theorem 5.10 $\mathfrak{g}_{M}$ does not admit a symplectic structure.
As a final application of Theorem 4.10 and Theorem 5.10, we obtain families of nilpotent almost abelian Lie algebras admitting both, complex and symplectic structures, or symplectic but no complex structure, or neither complex nor symplectic structure.